# Identifiability of VAR(1) model in a stationary setting

### v2

@@ Abstract
We consider a classical First-order Vector AutoRegressive (VAR(1)) model, where we interpret the autoregressive interaction matrix as influence relationships among the components of the VAR(1) process that can be encoded by a weighted directed graph. A majority of previous work studies the structural identifiability of the graph based on time series observations and therefore relies on dynamical information. In this work we assume that an equilibrium exists, and study instead the identifiability of the graph from the stationary distribution, meaning that we seek a way to reconstruct the influence graph underlying the dynamic network using only static information. Importantly, the graph associated to the support of the autoregressive interaction matrix of the VAR(1) model does not coincide with the graphical model of the stationary distribution, making standard tools developped for Gaussian Graphical Models (GGMs) irrelevent. Therefore, we use an approach from algebraic statistics that characterizes models using the Jacobian matroids associated with the parametrization of the models, and we introduce sufficient graphical conditions under which different graphs yield distinct steady-state distributions. Additionally, we illustrate how our results could be applied to characterize networks inspired by ecological research.
@@ 1 Introduction
In this paper, we focus on the identifiability of the underlying influence graph of the First-order Vector AutoRegressive (VAR(1)) model in a stationary setting. The VAR(1) model considered in this paper is:
$\left\{\begin{array}[]{l}\mathbf{x}_{0}=\bm{\epsilon}_{0},\\ \mathbf{x}_{t}=\Lambda^{T}\mathbf{x}_{t-1}+\bm{\epsilon}_{t}\quad\mbox{for }t\in\mathbb{N},\\ \end{array}\right.$ (1)
where $\mathbf{x}_{t}$ and $\bm{\epsilon}_{t}$ are random vectors in $\mathbb{R}^{n}$ for some $n\in\mathbb{N}$. $\bm{\epsilon}_{t}$ is independent of $\mathbf{x}_{t-1}$, and for all $t^{\prime}\neq t$, $\bm{\epsilon}_{t^{\prime}}$ is independent of $\bm{\epsilon}_{t}$. $\Lambda=\left(\lambda_{ij}\right)$ is a deterministic $n\times n$ matrix with values in $\mathbb{R}$, and in this paper, we call it the interaction matrix. Each element $\lambda_{ij}$ represents the direct influence of $\mathbf{x}_{i}$ on $\mathbf{x}_{j}$. In this paper, we assume that the error term is centered with covariance matrix $\omega I_{n}$, where $I_{n}\in M_{n}(\mathbb{R})$ is the identity matrix, and $\omega\in\mathbb{R}^{+}$ is a positive constant.
If the eigenvalues of $\Lambda$ are all smaller than $1$ in absolute value, then as $t$ goes to infinity, $x_{t}$ converges in distribution to
$\mathbf{x}_{\infty}\sim\mathcal{N}\left(0,\Sigma\right),$ (2)
where $\Sigma=\left(\sigma_{ij}\right)\in M_{n}(\mathbb{R})$ ([2]). $\Sigma$ satisfies the following equation ([18]):
@@
In fact, (3) is equivalent to the following equation ([18]):
@@
where vec is the vec operator ([12, Section 10.2]) and $\otimes$ is the matrix tensor product. Therefore, $\Sigma$ is unique given a unique set of parameters $\Lambda$ and $\omega$.
@@
This paper studies the identifiability of the VAR(1) model in this stationary setting, i.e., whether different parameters ($\Lambda$ and $\omega$) produce the same $\Sigma$, using (3). In particular, we restrict the problem to the identifiability of the support of $\Lambda$ (i.e. the underlying graph structure), that is, we seek a way to locate its nonzero elements. Importantly, conclusions from Gaussian Graphical Models (GGMs) (see [10] for an introduction) are inapplicable to our settings because the graphical model of the stationary distribution, i.e. the graph corresponding to the support of the inverse of the covariance matrix $\Sigma^{-1}$ is not the one that is induced by the support of $\Lambda$. For instance, Example 1 illustrates a case where $\Sigma^{-1}$, which is calculable from $\Sigma$, is complete (i.e. all elements are non-zero) while $\Lambda$ is not. As a consequence, it is necessary to seek for other methods to treat the problem.
The study of identifiability in a stationary setting is particularly promising for applications in ecological research. For example, the Lotka-Volterra equations (see [1]) used to describe the evolution of species abundances can be resumed to VAR(1) models (PhD thesis in progress [11]). Time-series observations are not always available as species might be only sampled once per site, and the identifiability of Lotka-Volterra models in this framework is still unclear.
To the best of our knowledge, the identifiability of VAR(1) parameters in a stationary setting has rarely been explored in the existing literature. A majority of previous work that has dealt with the identifiability of the VAR(1) model, and achieved significant and systematic results, has relied on observations that were sampled as time series (e.g. [6], [15]). This is reasonable because the parameters involve dynamic interactions among the different components of the random process $\mathbf{x}_{t}$, making it natural to expect that it should be inferred from dynamic data. Nevertheless, there is a literature that deals with the identifiability of the interaction matrix $\Lambda$ of VAR(1) in a stationary setting. [18] provides a necessary identifiability condition based on the number of parameters to estimate by solving directly the quadratic equation (3). However, this condition is generally too easily satisfied, meaning that it only excludes a small subset of cases where the interaction matrix is not identifiable, leaving the majority of cases unresolved.
@@
To be more precise about the graphical model we consider in this paper, the VAR(1) model (1) is naturally associated with a directed graph $G=\left(V,\mathfrak{E}_{G}\right)$ with node set $V=\left\{1,2,\cdots,n\right\}$, each corresponding to one of the components of $\mathbf{x}_{t}$, and edge set $\mathfrak{E}_{G}$ consists of ordered pairs $\left(i,j\right)$ or $i\rightarrow j$, which represents the direct influence among components of $\mathbf{x}_{t}$. For any $i,j\in\left[n\right]$, an edge $\left(i,j\right)$ in $\mathfrak{E}_{G}$ indicates $\lambda_{ij}\neq 0$, meaning that there is a direct effect of $\mathbf{x}_{i}$ on $\mathbf{x}_{j}$. Figure 1 shows an example of a graphical VAR(1) model with $4$ nodes, and the corresponding graph $G$, where self-loops are omitted. Because $\Lambda$ is deterministic, i.e. it does not change over time, the dynamic model in $(a)$ is often encoded by the graph $G$ in $(b)$ for simplicity. We assume throughout the paper that self-loops always exist in $G$, i.e. $(i,i)\in\mathfrak{E}_{G},\forall i\in\left[n\right]$. It is reasonable as it aligns with many real-life scenarios. For example, in ecological networks, we expect the evolution of species to be dependent on itself.
@@
We address the problem using the methods from algebraic statistics (see [14] for an introduction). Specifically, each graph is associated with a set of possible stationary distributions, and we represent them by the set of corresponding covariance matrices $\Sigma$, denoted as $\mathbf{M}_{G}$. We study the naturally associated algebraic variety defined as the Zariski closure of $\mathbf{M}_{G}$. Starting with a finite family of graphs, we study the intersections of the corresponding $\mathbf{M}_{G}$’s, which is particularly useful in cases where we can reduce the problem to a finite number of graph candidates. The algebraic approach used in this paper has already been proven fruitful by [4] in the context of linear structural equation models (SEMs), for which sufficient graphical conditions for the identifiability are derived.
We apply the methods for testing the identifiability of [4] to our settings and set up a framework for the problem of identifiability of the graphs of the VAR(1) model in a stationary setting. These methods characterize models using the Jacobian structures of the parametrizations. We introduce a new concept called “maximal classes” for directed graphs, which can be interpreted as sets of nodes that receive information from a common ”source” (detailed definitions in later sections). We present sufficient graphical conditions for identifiability based on maximal classes and give algorithms to reconstruct these maximal classes from any graph. In the end, we provide several illustrations of how to apply our results to networks inspired by ecological research.
@@
The rest of this paper is organized as follows: Section 2 formally defines the statistical model and introduces the definitions and testing criteria for global and generic identifiability. Additionally, we provide an overview of our main results of this paper in the end. Section 3 presents results on the Jacobian structure of the parametrization of the model. Section 4 presents a new concept called “maximal classes” for directed graphs, which is the core of this study. Section 5 introduces two sufficient conditions of generic identifiability based on maximal classes, one of which is only valid for characterizing models with the same dimension, while the other one can be extended to models with any dimension. Additionally, we present a result that allows us to calculate the dimension of the model for graphs without ”multi-edges”. Section 6 is an illustration of how our results could be applied in ecological research.
@@ 2 Identifiability of the VAR(1) model in a stationary setting
In this section, we start with introducing the parametrization and statistical model of our settings, then define global and generic identifiability, along with the identifiability criteria used throughout this paper. Finally, we provide an overview of our main results, i.e. sufficient conditions for generic identifiability.
@@ 2.1 Problem settings
where $\rho(\cdot)$ is the spectral radius, and the condition $\rho(\Lambda)<1$ ensures the existence of the equilibrium. The parameter space is then $M_{G}\times\mathbb{R}^{+}$, corresponding to elements of $\Lambda$ and $\omega$. Note that the parameter space includes the case where $\left(i,j\right)\in\mathfrak{E}_{G}$ and $\lambda_{ij}=0$ so that if the graphs $G_{1}=\left(V,\mathfrak{E}_{G_{1}}\right)$ and $G_{2}=\left(V,\mathfrak{E}_{G_{2}}\right)$ are nested, i.e. $\mathfrak{E}_{G_{1}}\subseteq\mathfrak{E}_{G_{2}}$, then the sets of possible stationary distributions generated from the two models are also nested.
@@
For a finite family of stationary VAR(1) models, we characterize them by studying the dimension of the intersection of the images of the parametrization map, in the sense that two models are generically identifiable if they only intersect in a set with smaller dimension (detailed definitions of identifiability introduced in the next section). Here, the dimension of $\mathbf{M}_{G}$ refers to the algebraic dimension with respect to the Zariski topology as an algebraic variety (see [14] for a full description). Because our model is parametric, and the parametrization map $\phi_{G}$ is rational (proof in Appendix A), the dimension of the model $\mathbf{M}_{G}$ equals the rank of the (transpose) of the Jacobian matrix of the parametrization map evaluated at a generic point ([14, Chapter 16, Th.16.1.7]). Here, a generic point of an algebraic set is a point in a general position, at which all generic properties are true. Therefore, for simplicity, we define the dimension of the model as the rank of the (transpose) of the Jacobian matrix. Note that in this case, the algebraic dimension of $\mathbf{M}_{G}$ coincides with the dimension of $\mathbf{M}_{G}$ viewed as a manifold.
@@ 2.2 Global and generic identifiability
In fact, there exist many tools to test generic identifiability [14, Chapter 16]. This paper applies one of them that uses the Jacobian matroid derived from the parametrization of the model, developed by [4]. The following section provides an introduction to this method.
@@ 2.3 Identifiability criteria
Now we introduce the Jacobian matroid, which is the main object that we use to characterize models. It is a matroid defined by the column independence of the Jacobian matrix. Relevant definitions are listed below.
@@
A matroid =$\mathcal{M}=\left\{E,\mathcal{I}\right\}$ is a pair where $E$ is a finite set and $\mathcal{I}$ is a set of subsets of $E$ that satisfies:
  • $\emptyset\in\mathcal{I}$;
  • If $I^{\prime}\subseteq I\in\mathcal{I}$, then $I^{\prime}\in\mathcal{I}$;
  • If $I_{1},I_{2}\in\mathcal{I}$ and $|I_{2}|>|I_{1}|$, then there exists $e\in I_{2}\backslash I_{1}$ such that $I_{1}\cup e\in\mathcal{I}$,
where ||$\cdot$|| ⋅ | represents the cardinality of a set.
Definition 8.
Let $A\in M_{m\times n}\left(\mathbb{R}\right)$ be a matrix, then the matroid defined by the column independence of $A$ is the pair $\left\{E,\mathcal{I}\right\}$, where $E=\{1,\cdots,n\}$, representing the $n$ columns of $A$, and
=$\mathcal{I}=\left\{S\subseteq E\mid A^{S}\mbox{ are linearly independent}\right\},$
@@
The fact that the column independence of a matrix defines a matroid is proved in [17, Theorem 27].
@@
then the matroid defined by the column independence of $A$ is $\left\{E,\mathcal{I}\right\}$, where $E=\left\{1,2,3,4\right\}$, representing the $4$ columns of $A$ respectively, and
={,$\mathcal{I}=\left\{\emptyset,\left\{1\right\},\left\{2\right\},\left\{3\right\},\left\{4\right\},\left\{1,2\right\},\left\{1,3\right\},\left\{2,3\right\},\left\{2,4\right\},\left\{3,4\right\},\left\{1,2,3\right\},\left\{2,3,4\right\}\right\}.$
@@
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model, then the Jacobian matroid of the model, denoted as $\mathcal{J}\left(\mathbf{M}_{G}\right)$, is the matroid defined by the column independence of the Jacobian matrix.
Next, we introduce the criteria for generic identifiability based on Jacobian matroids. The criterion is extensively introduced in [14] and [8]. It studies parametric algebraic statistical models for discrete random variables, meaning that there exists a rational map, i.e. the parametrization map:
$\phi:\Theta\rightarrow\Delta_{p-1}=\left\{s\in\mathbb{R}^{p}\mid\sum_{i=1}^{n}s_{i}=1,s_{i}\geq 0\mbox{ for all }i\right\},$
whose image is the model of interest. In particular, the stationary VAR(1) model falls into this category because the natural parameter space is finite-dimensional.
Proposition 2 ([8, Proposition 10]).
Let $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ be two irreducible algebraic models which sit inside the probability simplex $\Delta_{p-1}$. Without loss of generality assume ${\operatorname{dim}}\left(M_{1}\right)\geq{\operatorname{dim}}\left(M_{2}\right)$. If there exists a subset $S$ of the coordinates such that
Proposition 2 is applicable to our settings because our model is parametrized, hence irreducible([14]).
@@
This is a direct result of Proposition 2. ∎
@@
$S\in\mathcal{J}\left(\mathbf{M}_{k_{1}}\right)\backslash\mathcal{J}\left(\mathbf{M}_{k_{2}}\right)\mbox{ or }S\in\mathcal{J}\left(\mathbf{M}_{k_{2}}\right)\backslash\mathcal{J}\left(\mathbf{M}_{k_{1}}\right).$
@@
Propositions 3 and 4 will be the main tools used throughout this paper to test generic identifiability.

### v3

@@ Abstract
We consider a classical First-order Vector AutoRegressive (VAR(1)) model, where we interpret the autoregressive interaction matrix as influence relationships among the components of the VAR(1) process that can be encoded by a weighted directed graph. A majority of previous work studies the structural identifiability of the graph is based on time series observations and therefore relies on dynamical information. By constract, this work aims to incorporate the sampling insufficiency problem in ecological research, that is, due to limited resources, time-series observations are often unavailable and what we have in practice is independently, identically distributed data sampled from the stationary distribution, making classical time-series methods inapplicable. Within this framework, we assume that an equilibrium of VAR(1) exists, and study instead the identifiability of the graph from the stationary distribution, meaning that we seek a way to reconstruct the influence graph underlying the dynamic network using only static information. Importantly, the graph associated to the support of the autoregressive interaction matrix of the VAR(1) model does not coincide with the graphical model of the stationary distribution, making standard tools developped for Gaussian Graphical Models (GGMs) irrelevent. Therefore, this paper applies an approach from algebraic statistics that characterizes models using the Jacobian matroids associated with the parametrization of the models. We successfully derive easily applicable sufficient graphical conditions under which different graphs yield distinct steady-state distributions through the lens of maximal class, a concept that is initially introduced by this paper. Additionally, we illustrate how our results could be applied to characterize networks inspired by ecological research.
@@ 1 Introduction
VAR(1) model.
This paper considers a classical First-order Vector AutoRegressive (VAR(1)) model:
[…]
where $\mathbf{x}_{t}$ and $\bm{\epsilon}_{t}$ are random vectors in $\mathbb{R}^{n}$ for some $n\in\mathbb{N}$. $\bm{\epsilon}_{t}$ is independent of $\mathbf{x}_{t-1}$, and $(\bm{\epsilon}_{t})_{t\in\mathbb{N}}$ are i.i.d.. $\Lambda=\left(\lambda_{ij}\right)\in M_{n}\left(\mathbb{R}\right)$, i.e. it is a deterministic $n\times n$ matrix with values in $\mathbb{R}$, which we call the interaction matrix. Each element $\lambda_{ij}$ represents the direct influence of the $i^{\mbox{th}}$ coordinate of $\mathbf{x}_{t-1}$ on the $j^{\mbox{th}}$ coordinate of $\mathbf{x}_{t}$. In this paper, the error term $\bm{\epsilon}_{t}$ is centered with covariance matrix $\omega I_{n}$, where $I_{n}\in M_{n}(\mathbb{R})$ is the identity matrix, and $\omega\in\mathbb{R}^{+}$ is a positive constant.
If the eigenvalues of $\Lambda$ are all smaller than $1$ in absolute value, then as $t$ goes to infinity, $x_{t}$ converges in distribution to a centered distribution with covariance matrix $\Sigma=\left(\sigma_{ij}\right)\in M_{n}(\mathbb{R})$ ([2]). $\Sigma$ satisfies ([21]):
@@
Apply the vec operator ([15, Section 10.2]) on both sides, (2) is equivalent to:
@@
where $\otimes$ is the matrix tensor product ([21]). (3) indicates that $\Sigma$ is unique given a unique set of parameters $\Lambda$ and $\omega$.
Problem statement.
@@
This paper studies the identifiability of the VAR(1) model in a stationary setting, that is, whether different parameters ($\Lambda$ and $\omega$) produce the same $\Sigma$, using (2). We restrict the problem to the identifiability of the support of $\Lambda$, i.e. the location of its non-zero elements, which corresponds to the reconstruction of the underlying graph structure. Importantly, the properties of Gaussian Graphical Models (GGMs) (see [12] for an introduction) are inapplicable to our setting because the graphical model of the stationary distribution is the graph corresponding to the support of the inverse of the covariance matrix $\Sigma^{-1}$, and it is not the one induced by the support of $\Lambda$. For instance, Example 1 illustrates a case where $\Sigma^{-1}$, which is calculable from $\Sigma$, is complete (i.e. all elements are non-zero) while $\Lambda$ is not. Hence, it is necessary to seek for other methods to treat the problem.
@@
To be more precise about the graphical model, we consider in this paper, the VAR(1) model (1) is associated with a directed graph $G=\left(V,\mathfrak{E}_{G}\right)$ with node set $V=\left\{1,2,\cdots,n\right\}$, each corresponding to one of the components of $\mathbf{x}_{t}$, and edge set $\mathfrak{E}_{G}$ consists of ordered pairs $\left(i,j\right)$ or $i\rightarrow j$, which represents the direct influence among components of $\mathbf{x}_{t}$. For any $i,j\in\left[n\right]$, an edge $\left(i,j\right)$ in $\mathfrak{E}_{G}$ indicates $\lambda_{ij}\neq 0$, meaning that there is a direct effect of the $i^{\mbox{th}}$ coordinate of $\mathbf{x}_{t-1}$ on the $j^{\mbox{th}}$ coordinate of $\mathbf{x}_{t}$. Figure 1 shows an example of a graphical VAR(1) model with $4$ nodes, and the corresponding graph $G$, where self-loops are omitted. Because $\Lambda$ is deterministic, i.e. it does not change over time, the dynamic model in $(a)$ is encoded by the graph $G$ in $(b)$ for simplicity. We assume throughout the paper that self-loops always exist in $G$, i.e. $(i,i)\in\mathfrak{E}_{G},\forall i\in\left[n\right]$, which is reasonable as it aligns with many real-life scenarios such as ecological networks where we expect the evolution of species to be dependent on itself.
@@
Motivation.
The motivation of this paper can be found in ecological research. Ecosystems are often modeled by dynamical systems. For example, in the model of Robert May ([14]), the linearization of the dynamics of abundancies or densities of $n$ interacting species around their equilibrium leads to the system:
$\frac{\mathrm{d}\mathbf{x}}{\mathrm{d}t}=\mathbf{Ax},$ (4)
where $\mathbf{A}$ is a deterministic $n\times n$ matrix that represents interactions among species through various mechanisms such as competition, predation or facilitation. The equation (1) can be seen as a stochastic and discrete in time version of (4).
VAR(1) models can also arise when we consider the (vectorial) fluctuation process $\mathbf{x}_{t}$ associated with individual-based models converging to Lotka-Volterra ordinary differential equations. The fluctuation process is then shown to satisfy a continuous time version of (1) with a noise having a diagonal covariance matrix (see [1] for more precisions). More importantly, it is reasonable to start from a basic and classical model also because, as pointed out later in this section, the study of dynamical system with only information from a steady state is a subject rarely explored.
The reason for which we are particularly considering a stationary setting lies in the sampling insufficiency problem of ecological research. It is a central topic for ecologists to reconstruct the dynamic interaction network, such as the support of $\mathbf{A}$ in (4) or the support of $\Lambda$ in (1), from the abundance data captured in the ecosystem, since studying the dynamic influence among species is crucial to understanding the evolution of the abundances of species and thus helps to prevent extinction. However, as pointed out by [13], fine-scale replication sampling such as time-series sampling is generally too costly and are often sacrificed in practice. Instead, the limited resources (manpower, time, money) are used to cover as much ground as possible during each sampling campaign. Consequently, a considerable amount of abundance data available covers a large number of locations but each of them is sampled only once, making classical time-series analysis inapplicable. For example, [4] inferred a global ocean cross-domain plankton co-occurrence network with one-time-samplings of $200$ stations. It is true that there are cases where each of the sampling sites is sampled multiple times, but the temporal gap between each sampling campaign is generally too large. For instance, [22] sampled for plant–aphid–ant interactions on $18$ forest islands once a year from 2020 to 2021. There are also researches that collect data regularly, e.g. monthly, for each sampling site, but the data points are not enough to conduct classical time-series analysis, such as [17], which collected plant and humming bird abundance over $19$ sites monthly but only over a little more than a year.
To resolve the sampling insufficiency problem and potentially make the planning of sampling campaigns easier for ecologists, we seek for a way to reconstruct the dynamic interaction network of an ecosystem with only i.i.d. samples of a final stationary distribution of the dynamical system, instead of time-series observations. As a first step of solving this problem, this paper assumes that we have the final state of the VAR(1) model, i.e. the matrix $\Sigma$ in (2), from which we try to recover the dynamic interaction network, i.e. the support of $\Lambda$. Although in practice, we will have i.i.d. observations from the stationary distribution and never the true $\Sigma$, it is worth verifying that the network is identifiable with the true $\Sigma$, because if not, it’s highly unlikely to be identifiable with real data.
Related work.
To the best of our knowledge, the identifiability of VAR(1) parameters in a stationary setting has rarely been explored in the existing literature. A majority of previous work that has dealt with the identifiability of the VAR(1) model, and achieved significant and systematic results, relied on observations that were sampled as time series (e.g. [7], [19]). This is reasonable because the parameters involve dynamic interactions among the components of the random process $\mathbf{x}_{t}$, making it natural to expect that it should only be inferred from dynamic data. Nevertheless, there is a literature that deals with the identifiability of the interaction matrix $\Lambda$ of VAR(1) in a stationary setting. [21] provides a necessary identifiability condition based on the number of parameters to estimate by counting the parameters. However, this condition only excludes a small subset of cases where the interaction matrix is not identifiable, while we are more interested in sufficient conditions for identifiability. Young’s condition also didn’t provide enough insight into the link between dynamical network and the stationary distribution it generates, resulting in limited interpretability.
Our contribution.
We address the problem using methods from algebraic statistics (see [18] for an introduction). These algebraic approaches have already been proven fruitful by [5] in the context of linear structural equation models (SEMs), for which sufficient graphical conditions for identifiability are derived. We apply the methods for testing the identifiability of [5] to our settings and set up a framework for identifiability of VAR(1) graphs in a stationary setting. We represent the set of stationary distributions corresponding to a graph $G$ by the set of covariance matrices $\Sigma$, denoted as $\mathbf{M}_{G}$, and study the naturally associated algebraic variety defined as the Zariski closure of $\mathbf{M}_{G}$. In algebraic statistics, a finite family of graphs are generically identifiable (formal definition in Section 2) if the intersection of any pair of models $\mathbf{M}_{G_{1}}$, $\mathbf{M}_{G_{2}}$ associated with graphs $G_{1}$, $G_{2}$ is a Lebesgue measure zero subset of any of the two models, which could be characterized through the Jacobian structures of the parametrization. By applying these results into our setting, we successfully derive sufficient graphical conditions for generic identifiability. Our results are based on a concept that we introduced in this paper: “maximal classes”, which can be interpreted as sets of nodes that receive information from a common ”source” in a directed graph (formal definition and relevant reconstruction algorithms are provided in later sections). We prove that graphs with different maximal classes generally yield different stationary distributions. Our graphical conditions for generic identifiability are not only easy to verify in practice, but also highly interpretable, as maximal classes serve as a bridge between dynamical network and the corresponding stationary distribution. We also provide several illustrations of how to apply our results to networks inspired by ecological research.
Outline.
@@
The rest of this paper is organized as follows: Section 2 formally defines the statistical model and introduces the definitions and testing criteria for global and generic identifiability with an overview of the main results of this paper in the end. Section 3 presents results on the Jacobian structure of the parametrization of the model. Section 4 presents the concept of “maximal classes” for directed graphs and relevant reconstruction algorithms, which is the core of this study. Section 5 introduces two sufficient conditions of generic identifiability based on maximal classes, one of which is only valid for characterizing models with the same dimension, while the other one can be extended to models with any dimension. A result that allows us to calculate the dimension of the model for graphs without ”multi-edges” is also provided to make the two conditions applicable. Section 6 is an illustration of how our results could be applied in ecological research.
@@ 2 Identifiability of the VAR(1) model in a stationary setting
In this section, we start with introducing the parametrization and statistical model of our settings, then define global and generic identifiability, along with the identifiability criteria used throughout this paper. Finally, we provide an overview of our main results, i.e., sufficient conditions for generic identifiability.
The definition and criteria for identifiability are based on the concept of a generic point. We define a point in a set as generic if it does not lie within a proper (i.e., strictly included) algebraic subvariety of the set. In concrete terms, a generic point is any point in the set except for those that satisfy a specific non-trivial polynomial equation. Because such proper algebraic subvarieties have Lebesgue measure zero, another interpretation is that a parameter value sampled from a continuous probability distribution is generic with probability one.
@@ 2.1 Problem settings
where $\rho(\cdot)$ is the spectral radius, and the condition $\rho(\Lambda)<1$ ensures the existence of the equilibrium. The parameter space is then $M_{G}\times\mathbb{R}^{+}$, corresponding to elements of $\Lambda$ and $\omega$.
@@
Remark 1 (Structural Properties).
We highlight two structural characteristics of the parameter space $M_{G}\times\mathbb{R}^{+}$ and the model $\mathbf{M}_{G}$:
1. Nested structure: The definition of $M_{G}$ allows for edge weights to vanish (i.e., $\lambda_{ij}=0$ for $(i,j)\in\mathfrak{E}_{G}$). Consequently, the parameter space of any subgraph $G^{\prime}\subset G$ is naturally embedded within $M_{G}$. This inclusion property extends to the models (i.e., $\mathbf{M}_{G^{\prime}}\subseteq\mathbf{M}_{G}$), ensuring a continuous transition between the model $\mathbf{M}_{G}$ and its sub-models.
2. Error structure: The covariance matrix of the error term in (1) is assumed to be diagonal with identical values $\omega$. While this assumption is central to our calculations, the underlying framework is generalizable to other diagonal matrices with distinct entries. However, the calculations do not extend to non-diagonal matrices. We adopted the identical diagonal assumption to minimize the number of parameters to estimate, allowing us to focus specifically on the reconstruction of $\Lambda$.
@@
For a finite family of stationary VAR(1) models, we characterize them by studying the dimension of the intersection of the images of the parametrization map, in the sense that two models are generically identifiable if they only intersect in a set with smaller dimension (detailed definitions of identifiability introduced in the next section). The dimension of the model $\mathbf{M}_{G}$ is defined as the rank of the (transpose of the) Jacobian matrix of the parametrization map evaluated at a generic point. This definition coincides with the standard dimension of $\mathbf{M}_{G}$ viewed as a manifold. Formally, this corresponds to the algebraic dimension with respect to the Zariski topology as an algebraic variety (see Appendix A for details).
@@ 2.2 Global and generic identifiability
In fact, there exist many tools to test generic identifiability [18, Chapter 16]. This paper applies one of them that uses the Jacobian matroid derived from the parametrization of the model, developed by [5]. The following section provides an introduction to this method.
@@ 2.3 Identifiability criteria
Now we introduce the Jacobian matroid, the main object used to characterize models. To do so, we first introduce the linear matroid associated with a matrix. The general definition of a matroid is omitted here, as it is not necessary for the tools used in this paper. We refer the interested readers to [18] for further details.
@@
Let $A\in M_{m\times n}\left(\mathbb{R}\right)$ be a matrix, then the linear matroid of $A$ is the pair $\left\{E,\mathcal{I}\right\}$, where $E=\{1,\cdots,n\}$, representing the $n$ columns of $A$, and
$\mathcal{I}=\left\{S\subseteq E\mid A^{S}\mbox{ are linearly independent}\right\},$
@@
then the linear matroid of $A$ is $\left\{E,\mathcal{I}\right\}$, where $E=\left\{1,2,3,4\right\}$, representing the $4$ columns of $A$ respectively, and
$\mathcal{I}=\left\{\emptyset,\left\{1\right\},\left\{2\right\},\left\{3\right\},\left\{4\right\},\left\{1,2\right\},\left\{1,3\right\},\left\{2,3\right\},\left\{2,4\right\},\left\{3,4\right\},\left\{1,2,3\right\},\left\{2,3,4\right\}\right\}.$
@@
The Jacobian matroid of a stationary VAR(1) model $\mathbf{M}_{G}$, denoted as $\mathcal{J}\left(\mathbf{M}_{G}\right)$, is the linear matroid of the Jacobian matrix.
By abuse of notation, we identify $\mathcal{J}\left(\mathbf{M}_{G}\right)$ with $\mathcal{I}$ in Definition 7, representing the set of linearly independent columns sets of the Jacobian matrix.
Next, we introduce the criteria for generic identifiability based on Jacobian matroids. The criterion is extensively introduced in [18] and [10].
@@
We refer the reader to [10, Proposition 10] for a proof of this result. ∎
@@
[…]
@@
Propositions 2 and 3 will be the main tools used throughout this paper to test generic identifiability.